OpenStudy (anonymous):

find out this:f is a quadratic function whose graph is a parabola opening upward and has a vertex on the x-axis. The graph of the new function g defined by g(x) = 2 - f(x - 5) has a range defined by the interval A. [ -5 , + infinity) B. [ 2 , + infinity) C. ( - infinity , 2] D. ( - infinity , 0]

OpenStudy (dumbcow):

f(x) = a(x-h)^2 f(x-5) = a(x-h-5)^2 g(x) = 2-a(x-h-5)^2 g is a parabola which opens down and has bee shifted 2 units up range is (-inf, 2]